OpenStudy (anonymous):

Find the GCF and LCM of the following monomials 28,35 and 42 plz help!!

OpenStudy (anonymous):

7 and 420 i think

OpenStudy (cwrw238):

write each number in prime factor form 28 = 2*2*7 35 = 5*7 42 = 2*3* 7 7 is the only factor common to all 3 so thats the GCF LCM; choose one of the 3 7's, one 2 of two 2's, and 2,3 and 5 = 2*2*3*5*7 = 12*5*7 = 420 alpha is rght

OpenStudy (anonymous):

Thx i was 95%

OpenStudy (cwrw238):

no probs

OpenStudy (anonymous):

I am always 95 idk y

OpenStudy (cwrw238):

lol - try to be more confident!